OpenStudy (anonymous):

solve equation and please show work: 2(x+6)/3=x-5

OpenStudy (anonymous):

2(x+6)/3=x-5 (2x+12)/3=x-5 2x+12=3x-15 27=x

OpenStudy (anonymous):

medal??? :)

OpenStudy (anonymous):

Thanks!